1.   A tightly wound 1500-turn toroid has an inner radius 1.00 cm and an outer radius 2.00 cm, and carries a current of 2.50A. What is the magnetic field strength at a distance of 1.20 cm from the origin? 



1
Expert's answer
2021-11-24T11:59:59-0500

Inner radius a = 1.00 cm

Outer radius b = 2.00 cm

I = 2.50 A

N = 1500

r=1.20

"a<r<b \\

B = \frac{\mu_0 N I}{2 \pi r} \\

B = \frac{4 \pi \times 10^{-7} \times 1500 \times 2.5}{2 \pi \times 0.012} \\

B = 0.0625 = 62.5 \;mT"